Hello,
I have a question.
An actor invest some money at 5% and $27000 more than twice the amount at 11%. The total annual interest earned from the investment is $34560. How much did he invest at each amount.
Thank you

Answer by mananth(16949)

Amount @ 5.00% x
Amount @ 11.00% 2 x + 27000
Total interest earned = 34560

5.00% x+ 11.00% ( 2 x + 27000 )= 34560
5.00% x+ 11.00% * 2 x + 2970 = 34560
5.00% x+ 0.22 x + 2970 = 34560
0.05 x+ 0.22 x = 31590
0.27 x = 31590
x= 117000 Amount @ 5.00%

Amount @ 11.00% = 2 * 117000 + 27000
= 261000